Fermi National Laboratory
 

Accelerator Update for January 16, 2009 - January 19, 2009

 

Friday Jan 16

The day shift began with Operations monitoring the Tevatron (TeV) store 6730, the Recycler stash (368.2E10), and the Antiproton Source stack (24.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

Operations terminated store 6730 at 8:15 AM.

Operations put the TeV into shot setup at 8:29 AM.

At 8:44 AM, two MI kickers (I:KPS4A & B) failed to fire for the first of four times in about an hour. Kicker experts investigated.

Operations established store 6731 at 10:18 AM with an initial luminosity of 308.83E30.

At 10:32 AM, a Preaccelerator expert accessed the I- Source dome.

Linac experts began replacing the power amplifier for Linac RF station #1 (LRF1) at 10:33 AM.

Linac experts began work on LRF2 at 10:36 AM.

A Booster expert reported at 11:23 AM that he was reterminating a kicker cable (MKS04).

A Recycler expert began a study at 11:40 AM.

Linac experts reported at 11:47 AM that they had worked on many of the Klystron modulators.

At 3:26 PM, BRF11 tripped off on a ground fault indication. An RF expert investigated, bypassed the station, and then brought BRF19 online.

Operations established beam to the Linac dump at 3:51 PM.

Operations reported at 3:58 PM that the Recycler had stashed antiprotons twice during the day shift.

Operations established beam to Pbar at 4:09 PM.

At 5 PM, LRF2 tripped off and would not reset. An expert investigated.

Operations put Pbar into standby mode at 5:08 PM.

Linac experts reported that they had brought LRF2 back online at 5:44 PM.

At 7:07 PM, the Booster East Anode Power Supply tripped off and would not reset. Experts investigated and brought the power supply back online at 7:32 PM.

At 7:33 PM, Operations reported that BRF3 tripped off and that many other RF stations were sparking. Experts investigated and said the problem was in Linac.

Pbar resumed stacking at 8:31 PM.

At 10:46 PM, BRF6 tripped off. It reset.

At 11:32 PM, BRF8 tripped off. It reset.

Operations reported at 11:52 PM that the Recycler had stashed five times during the evening shift.

Saturday Jan 17

The midnight shift began with Operations monitoring store 6731 (81.65E30) and the stash (181.3E10), with Pbar stacking (25.6E10), and with NuMI and MiniBooNE taking beam.

At 12:55 AM, BRF6 tripped off several times. It reset, but then stations BRF7, 11, and 16 all started sparking.

Operations reported at 1:17 AM that the Booster RF sparking was due to Linac. Operators adjusted the KRF7 phase and centered the Linac velocity meter and the sparking stopped.

At 5:23 AM, BRF8 tripped off. It reset.

At 6:13 AM, the MI South Anode Power Supply tripped off. It reset.

Operations terminated store 6731 at 7:08 AM.

Operations prepared D0 and the TeV sectors A-E at 7:49 AM for access.

Operations reported at 8:02 AM that the Recycler had stashed nine times during the midnight shift.

At 8:26 AM, a Linac quadrupole power supply (L:QPS219) inhibited beam due to being near its high current limit.

At 8:37 AM, a MI kicker (I:KOS4A) tripped on a prefire fault.

At 10:05 AM, Operations reported that the Recycler was having cooling problems.

D0 came out of access at 10:33 AM.

TeV came out of access at 10:51 AM.

At 11:04 AM, BRF6, 7, & 8 began tripping off and reset.

At 11:39 AM, two MI kickers tripped off on a prefire fault.

Operations reported at 12:23 PM that experts had repaired the Recycler cooling problems.

Operations reported at 12:33 PM that experts discovered a guardrail, which had been removed to make repairs, at TeV sector E11 leaning against a bus.

At 1:01 PM, Operations reported that many Booster RF stations were sparking.

At 1:56 PM, Operations reported that the TeV sector A3 VCB would not close in. Experts responded.

Linac experts again made adjustments to KRF7 and D4VELO at 2:07 PM to resolve the Booster RF trips.

TeV power supply experts reported at 3:23 PM that they had closed in the TeV sector A3 VCB switch.

At 3:58 PM, KRF5 tripped off. It reset, but then a number of Booster RF stations tripped off.

An expert reported at 5:36 PM that the problem with the Linac Velocity Meter was due to a diagnostic scope that had been connected to monitor system signals. It was disconnected.

Operations put the TeV through a wet squeeze at 5:51 PM

Operations put the TeV into shot setup at 6:58 PM.

At 7:22 PM, Operations reported a plate-to-plate spark on the TeV sector C17 vertical separator. Experts reported that the spark was due to calculation errors and that the beam was fine.

Operations established store 6735 at 8:39 PM with an initial luminosity of 330.6E30.

At 10:14 PM, KRF1 tripped off on a LCW flow indication. It reset.

At 10:55 PM, a TeV RF station (TRF6) tripped off. This trip caused a small loss of luminosity to store 6735.

At 11:54 PM, two MI kickers (KPS4A & B) tripped off on a prefire fault for the first of five times in about four hours. It reset.

Sunday Jan 18

The midnight shift began with Operations monitoring store 6735 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 12:38 AM, an operator reported that he had switched a Linac quadrupole power supply to its shunt mode.

Operations reported at 8:01 AM that the Recycler had stashed 12 times during the last 12 hours.

At 8:14 AM, BRF6, 7, 8 and 16 tripped off. They reset.

At 8:40 AM, KRF5 tripped off. It reset.

Operations terminated store 6735 at 9:04 AM.

Operations put the TeV into shot setup at 9:17 AM.

An RF expert reported at 9:18 AM that he was turning on TRF6.

Operations established store 6736 at 10:50 AM with an initial luminosity of 339.89E30.

At noon, KRF4 tripped off. It reset.

At 12:01 PM, BRF8 tripped off. It reset.

At 2:02 PM, BRF3, 4, and 7 tripped off. They reset.

At 2:33 PM, BRF8 tripped off. It reset.

At 2:34 PM, LRF3 tripped off. It reset.

At 2:45 PM, BRF15 and 16 tripped off after a vacuum pump tripped off (IPRF24). They reset.

At 4:07 PM, the NuMI beam permit tripped off on MI RF incomplete status, but no RF stations were down. Everything reset.

At 5:04 PM, BRF3, 4, and 7 tripped off. They reset.

At 5:44 PM, two MI kickers (KPS4A and B) tripped off on a prefire fault. They reset.

At 6:46 PM, a MI kicker (KPS5S) tripped off and caused the NuMI beam permit to also trip. Everything reset.

At 7:01 PM, BRF8 tripped off. It reset.

At 7:38 PM, BRF3, 4, and 7 tripped off. They reset.

At 7:40 PM, KPS5S tripped off and caused NuMI to trip off. It reset.

At 7:55 PM, LRF1 tripped off. It reset.

Operations reported at 7:59 PM that the Recycler had stashed 11 times during the last 12 hours.

Monday Jan 19

The midnight shift began with Operations monitoring store 6736 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 1:12 AM, the P1 transfer beam line beam permit tripped off due to a power supply’s (I:V714) output current drifting too high. Everything reset.

Operations terminated store 6736 around 2:29 AM.

Operations put the TeV into shot setup at 2:39 AM.

Operations established store 6737 at 4:01 AM with an initial luminosity of 349.2E30.

At 5:30 AM, a Pbar Debuncher power supply (D:IB) tripped off on a ground fault. Operations called in an expert.

Operations reported at 5:39 AM that D:IB had tripped off due a LCW leak in the power supply.

At 5:53 AM, Operations reported a TeV quench; store 6737 was lost.

Operations prepared CDF and D0 for an access at 7:07 AM.

A Pbar power supply expert reported at 7:16 AM that he had replaced a bad hose.

Operations reported at 8:02 AM that the Recycler had stashed seven times during the last 12 hours.

The Plans for Monday

The plans for today are to allow CDF and D0 to complete their access and to repair the Tevatron Electron Lens.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/20/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory